OpenStudy (anonymous):

How do I evaluate in the real number system: (^3sqr.root:5^2)(^3sqr.root:5)?

OpenStudy (anonymous):

\[(^{3}\sqrt{5}^{2})(^{3}\sqrt{5})\]

OpenStudy (anonymous):

\[\large (\sqrt[3]5^2)(\sqrt[3]5)=5^{\frac{2}{3}}\times 5^{\frac{1}{3}}=5^{\frac{2}{3}+\frac{1}{3}}=5^1=5 \]

OpenStudy (anonymous):

ohhh.. thank you!